OpenStudy (anonymous):

which of the flowing are derrivatives of f(x)=2/secxcscx. 1) sin^2 x+c 2)-cos^2 x+c 3)-cos2x/2 + c

OpenStudy (anonymous):

anti derivatives sorry

OpenStudy (anonymous):

\[\int \frac{2}{\sec x \csc x}dx=\int2 \sin x \cos x dx=\int\sin 2x dx=-\frac{1}{2}\cos 2x+C\]

OpenStudy (anonymous):

so the first one and last one ??

OpenStudy (anonymous):

or just last one ?

OpenStudy (anonymous):

Yes. Both of them, since \(\cos 2x=1-2\sin^2 x\).

OpenStudy (anonymous):

so the second one no right?

OpenStudy (anonymous):

Yep. :)

OpenStudy (anonymous):

thank you Inlucero <3

OpenStudy (anonymous):

Your welcome. :)